### v3

# REDUCIO! Generating 1K Video within 16 Seconds using Extremely Compressed Motion Latents

Abstract
Commercial video generation models have exhibited realistic, high-fidelity results but are still restricted to limited access. One crucial obstacle for large-scale applications is the expensive training and inference cost. In this paper, we argue that videos contain significantly more redundant information than images, allowing them to be encoded with very few motion latents. Towards this goal, we design an image-conditioned VAE that projects videos into extremely compressed latent space and decode them based on content images. This magic Reducio charm enables 64 reduction of latents compared to a common 2D VAE, without sacrificing the quality. Building upon Reducio-VAE, we can train diffusion models for high-resolution video generation efficiently. Specifically, we adopt a two-stage generation paradigm, first generating a condition image via text-to-image generation, followed by text-image-to-video generation with the proposed Reducio-DiT. Extensive experiments show that our model achieves strong performance in evaluation. More importantly, our method significantly boosts the training and inference efficiency of video LDMs. Reducio-DiT is trained in just 3.2K A100 GPU hours in total and can generate a 16-frame 10241024 video clip within 15.5 seconds on a single A100 GPU. Code is available at https://github.com/microsoft/Reducio-VAE.
Impressionist style, a yellow rubber duck floating on the waves at sunset.
An aerial shot of a lighthouse standing tall on a rocky cliff, its beacon cutting through the early dawn, waves crashing against the rocks below.
A wizard waves his wand.
1 Introduction
Recent advances in latent diffusion models (LDMs) [45] for video generation have presented inspiring results [71, 4, 1, 3, 61, 74, 2, 5, 41, 67, 56, 53, 55, 54], showing great potential in various applications. Commercial models like Sora [36], Runway Gen-3 [46], Movie Gen [40], and Kling [24] can already generate photorealistic and high-resolution video clips. However, training and deploying such models are computationally demanding—thousands of GPUs and millions of GPU hours are required for training, and the inference for a one-second clip costs several minutes. Such a high cost is becoming a significant obstacle for research and large-scale real-world applications.
Extensive effort has been made to relieve the computational burden by means of efficient computation modules in the backbone [52, 70], optimizing diffusion training strategies [14, 66], and adopting few-step sampling methods [49, 73]. However, in this paper, we argue that the very nature of the problem is long neglected. Concretely, most video LDMs stick to the paradigm of text-to-image diffusion, e.g., Stable Diffusion (SD) [45], and inherently employ the latent space of its pre-trained 2D variational autoencoder (VAE) [23, 43]. This scheme compresses the input videos by in each spatial dimension, which is suitable for images while being much surplus for video.
As videos naturally carry much more redundancy than images, we believe they can be projected into a much more compressed latent space with a VAE based on 3D convolutions. Surprisingly, for videos, a spatial down-sampling factor does not sacrifice much of the reconstruction performance. While the resulting distribution of latent space shifts from those of the pre-trained image LDMs, we can still wield their well-learned spatial prior by factorizing text-to-video generation with a condition frame [13, 63, 69, 78]. In particular, we can obtain a content frame with off-the-shelf text-to-image LDMs and then generate samples with text and image as joint priors into video LDMs.
More importantly, as the image prior delivers enriched spatial information of video contents, factorized video generation offers the potential for further compression. We argue that video can be encoded into very few latents that represent motion variables, i.e., motion latents, plus a content image. Hereby, we introduce the magical Reducio charm by building a VAE comprised of a 3D encoder that aggressively compresses input videos into a down-sampled compact space, and a 3D decoder fused with feature pyramids of the middle frame as the content condition. Notably, the resulting Reducio-VAE surpasses the common 2D VAE by 5db in PSNR while using a smaller latent space.
Subsequently, we establish the LDM with a diffusion transformer (DiT) [37, 6]. Besides using T5 features [42] as the text condition, we employ an image semantic encoder as well as a context encoder to provide additional image conditions, which inform the model of the spatial content. Thanks to the extremely compressed video latent, our diffusion model, namely Reducio-DiT, thus enjoys both the fast training (inference) speed and high generation quality. Specifically, Reducio-DiT achieves 318.5 FVD score on UCF-101, surpassing a bunch of previous work [4, 1, 64, 79, 63] while being much faster. It can also be easily upscaled to a larger resolution, e.g. , with affordable cost. Our experiments show that Reducio-DiT can generate a 16-frame video clip of resolution within 15.5 seconds on a single A100 GPU, and hence achieves a speedup over Lavie [64].
In summary, our contributions are as follows: (1) We encode video into very few motion latents plus a content image, where the designed Reducio-VAE can compress video into a smaller latent space compared to common 2D VAE. (2) Based on Reducio-VAE, a diffusion model called Reducio-DiT is devised by incorporating multimodal conditions to perform text-image to video generation. (3) Experiments show that our Reducio-DiT can significantly speed up the generation process while producing high-quality videos.
2 Related Work
Video diffusion models have become the latest craze in the field of text-to-video generation. Due to the huge success of text-to-image LDMs [17, 34, 48, 33, 45], attempts have been made to leverage the powerful UNet-based stable diffusion models pre-trained in 2D space by inserting temporal convolution and attention layers [12, 70, 4, 3, 61]. Recently, the strong performance of transformers has continued in image generation domain [60, 31, 10, 6, 25]. Consequently, recent works have focused on extending 2D diffusion transformer (DiT) [37] to spatiotemporal domain.
While the mainstream exploration [40, 85, 26] follows the paradigm of image diffusion models by using text as the only condition, an emerging stream of works [13, 63, 78] decompose text-to-video diffusion into two steps, i.e., text-to-image and text-image-to-video generation. The factorized framework encourages models to focus on motion modeling instead of challenging spatiotemporal joint modeling. In this work, we build an image-to-video model upon DiT. Thanks to the delicate design of Reducio-VAE, our diffusion model preserves fine details of image priors in generated videos and hence can be used for more high-demanding real-world applications, e.g., advertisement generation.
Efficient diffusion models have gained increasing attention due to the tremendous computational resources and time required for training and inferring samples, especially for high-resolution images and videos. On one hand, researchers have made attempts to build LDMs with more efficient architectures [11, 52, 68, 84]. For example, DIM [52] uses Mamba as its backbone to ease the computational burden compared to vanilla transformers. Another line of work leverages efficient diffusion techniques to perform fewer sampling steps during inference [28, 30, 49, 73, 80] or to accelerate convergence during training [14, 66]. For instance, CausVid [75] applies distribution matching distillation between a causal student video generation model and a bidirectional teacher model.
Our work focuses on designing a more compact latent space [38], which straightforwardly and effectively reduces the overall training and inference costs of LDM. PVDM [76] and CMD [77] propose projecting 3D video samples into three separate yet compact 2D latent spaces. However, this approach requires training separate diffusion models for each space, which inevitably leads to discrepancies during inference. Our Reducio-VAE takes after LaMD [19] to adopt a factorized, compact motion latent space with the help of a content frame for reconstruction. In contrast, we emphasize aggressive compression of the spatial dimension and suggest that a moderate temporal down-sampling factor provides the best trade-off between performance and efficiency. Concurrent research in the image domain also indicates that using a high spatial down-sampling factor on redundant visual content [51, 68], such as high-resolution images, significantly improves efficiency. We propose that it is crucial to recognize the importance of spatial redundancy in the video latent space.
3 Reducio
3.1 Reducio Video Autoencoder
Variational autoencoder (VAE) [23] projects input images or videos into a compressed latent space obeying a certain distribution and maps the latent sampled from the obtained distribution into RGB space with a decoder. Latent diffusion models [45] take advantage of the down-sampled latent space of VAE to support generation with improved efficiency. In this paper, we follow the common practice and build our video autoencoder upon VAE.
We denote the temporal down-sampling factor of a VAE as and the spatial one as . Given an input video with the shape of , the compressed latent as the input into diffusion models fall into the space of , where denotes the latent channel. Thanks to the superior performance of open-source Stable Diffusion (SD) [45] models, it has been a common practice to leverage SD-VAE to process images and videos into compressed latent. Typically, as shown in (a) of Fig. 2, SD-VAE is built upon 2D convolution with and .
Reducio charm on videos. As adjacent frames share large similarities in pixel space, video signals inherently contain more redundancy than images. Therefore, we apply a more aggressive down-sampling strategy for video autoencoder. Specifically, we employ 3D convolutions and increase the number of autoencoder blocks, increasing and to 4 and 16, as shown in (b) of Fig. 2, respectively. Though the overall down-sampling factor increases by , the compressed latent space suffices to achieve comparable pixel-space reconstruction quality, highlighting the potential for extremely shrinking video latent dimensions.
As for image-to-video generation, we argue that video latent can be further compressed as the content image contains rich visual details and substantial overlapping information. In such a scenario, video latents can be further compressed to represent the motion information. We can find similar spirits in the video codec [44], where carefully designed modules use reference frames as inputs to perform high-fidelity reconstruction of a given compressed code. Specifically, we propose to cast the Reducio charm on video latent space, as shown in (c) of Fig. 2. Generally, we choose the middle frame as the content guidance, compress input videos by an extreme down-sampling factor with a 3D encoder, and reconstruct from the latent space with the aid of content frame features pyramids. Specifically, we infuse content features of sizes varying in and via cross-attention within the decoder, adopt an extreme compression ratio, i.e., and . While Reducio-VAE projects videos into more compressed latent space than one of SDXL-VAE [39], it significantly outperforms the latter in reconstruction performance.
Note that we train Reducio-VAE on videos of resolution. To encode and decode high-resolution videos without exceeding the memory limit, we split the input video into overlapping tiles along the spatial dimension, following the approach used in Movie Gen [40]. In practice, we use a tile size of 256 (8 in latent space) with an overlap of 64 pixels (2 latents). Overlapping latents and decoded video regions are blended using linear combinations.
3.2 Reducio Diffusion Transformer
Thanks to the magical compression capability of Reducio-VAE, we project input videos into compressed latent space with a down-sampling factor of 4096, hence significantly speeding up the training and inference of diffusion models. Specifically, we adopt the DiT-XL [37] model and follow most of the design choices adopted by PixArt- [6], i.e., AdaLN-single modules, cross attention layers with Flan-T5-XXL [42] text conditions.
To adapt the image diffusion model to video, we consider two options: (1) Directly transforming the 2D attention to full 3D attention without adding additional parameters [74, 8]. (2) Adding temporal layers and transforming the model to perform 2D spatial attention plus 1D temporal attention [85, 78, 26]. We adopt option (1) by default. Studies on the evaluation of these two options are shown in Sec. 4.3.2. In addition, we introduce Reducio-DiT (illustrated in Fig. 3) with additional image-condition modules, as described below.
Content frame modules consist of a semantic encoder built upon pretrained OpenCLIP ViT-H [9] and a content encoder initialized with the SD2.1-VAE [45]. The former encoder projects content frames into a high-level semantic space while the latter one mainly focuses on extracting the spatial information. We concatenate the obtained features with text tokens output by T5 to form the image-text-joint condition, and then cross-attend them with the noisy video latent.
| Model | Downsample | Pexels | UCF-101 | ||||
| Factor | PSNR | SSIM | LPIPS | rFVD | rFVD | ||
| SD2.1-VAE [45] | 4 | 29.23 | 0.82 | 0.09 | 25.96 | 21.00 | |
| SDXL-VAE [39] | 4 | 30.54 | 0.85 | 0.08 | 19.87 | 23.68 | |
| OmniTokenizer [59] | 8 | 27.11 | 0.89 | 0.07 | 23.88 | 30.52 | |
| OpenSora-1.2 [85] | 16 | 30.72 | 0.85 | 0.11 | 60.88 | 67.52 | |
| Cosmos-VAE [35] | 16 | 30.84 | 0.74 | 0.12 | 29.44 | 22.06 | |
| 16 | 28.14 | 0.65 | 0.18 | 77.87 | 119.37 | ||
| Reducio-VAE | 16 | 35.88 | 0.94 | 0.05 | 17.88 | 65.17 | |
Scaling up to high-resolution videos consumes tremendous computational resources for vanilla video diffusion models. Nevertheless, Reducio-DiT relieves this limit to a great extent. To support high-resolution (e.g., ) video generation, we employ a progressive training strategy: 1) in the first stage of training, the model learns to align the video latent space with text-image prior by taking in a vast number of videos as input; 2) in the second stage, we fine-tune the model on videos with higher resolution, i.e., . In consequence, the content encoder augments more tokens with the noisy latent; 3) in the third stage, we conduct further fine-tuning on videos with spatial resolution around with multi-aspect augmentation.
To better accommodate the model to videos with varying aspect ratios, we inject the aspect ratio and size of input videos as embeddings [6] into the DiT model. We then add the dynamic size embedding with timestep embedding before being fed into the MLP. More importantly, for models in the first and the second stages, the content encoder cross-attends tokens with the noisy latent. As computational costs grow rapidly when resolution increases, we introduce a strategy to ensure the efficiency of Reducio-DiT on high-resolution videos. Specifically, we aim to shrink the content embedding by in the spatial dimension in each cross-attention layer. Inspired by Deepstack [32], we divide input tokens into 4 groups of grids and iteratively concatenate each group of tokens with the other condition tokens.
4 Experiments
4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied by a short text description. We train Reducio-VAE from scratch on videos with 16 FPS. To perform the VAE comparison, PSNR, SSIM [65], LPIPS [81] and reconstruction FVD (rFVD) [57] are employed as the evaluation metrics.
Reducio-DiT is trained on the above Pexels dataset and an internal dataset with 5M videos containing high-resolution text-video pairs. As described in Sec. 3.2, we adopt a multi-stage training strategy to train the model from low resolution to high resolution. We first train Reducio-DiT-256 using a batch size of 512 on 4 Nvidia A100 80G GPUs for around 900 A100 hours and initialize model weights with PixArt--256 [6]. Then we fine-tune on videos shortly for 300 A100 hours to obtain Reducio-DiT-512. In the third stage, we randomly sample video batches from 40 aspect ratio buckets, which is the same as the setting in PixArt-. We leverage 8 AMD MI300 GPUs to support a training batch size of 768 and fine-tune Reducio-DiT-1024 for 1000 GPU hours.
| Model | FVD | |
| UCF-101 | MSR-VTT | |
| VideoComposer [62] | 576.8 | 377.3 |
| I2VGen-XL [83] | 571.1 | - |
| DynamiCrafter [69] | 429.2 | 234.7 |
| Reducio-DiT | 318.5 | 291.9 |
We adopt DPM-Solver++ [29] as the sampling algorithm for efficient inference and set the sampling step to 20. For evaluation in Tab. 3, we use PixArt--256 and PixArt--1024 to generate the content frame for and videos, respectively. We denote out-of-memory as OOM and calculate the speed of generating 16-frame video clips on a single A100 80G GPU, including both text-to-image and text-image to video cost. For Reducio-DiT, We report the FVD and IS scores of Reducio-DiT-512 under the zero-shot setting on UCF-101 [50] and MSR-VTT [72]. Since our model performs image-to-video generation and requires a condition image prior for the text-to-video pipeline, we follow MicroCinema [63] to use SDXL [39] to generate the condition image and LLaVA-1.5 [27] to generate captions for FVD evaluations on UCF-101 and MSR-VTT. Additionally, we evaluate Reducio-DiT-512 on the recent video generation benchmark VBench [20].
| Model | Params | VBench | FVD | Throughput (#videos/sec) | GPU Hours | ||
| UCF-101 | MSR-VTT | A100 Training | |||||
| Make-A-Video [47] | - | - | 367.23 | - | - | - | - |
| CogVideo [18] | 7.8B | 67.01 | 701.59 | 1294 | 0.0023 | OOM | 368K |
| Show-1 [79] | 1.7B | 78.93 | 394.46 | 538.0 | 0.0016 | OOM | 56K |
| Lavie [64] | 1B | 77.08 | 526.30 | - | 0.1838 | 0.0039 | - |
| MicroCinema [63] | 2B | - | 342.86 | 377.4 | 0.0234 | OOM | 6K |
| Lumiere [2] | 2B | - | 332.49 | - | - | - | - |
| Reducio-DiT | 1.2B | 81.39 | 318.50 | 291.9 | 0.9824 | 0.0650 | 3.2K |
4.2 Main Results
Reducio-VAE preserves fine details of video inputs. Fig. 4 displays the first and the last frame of videos reconstructed by SDXL-VAE and Reducio-VAE, with the yellow bounding box highlighting the region with obvious differences in details. While SDXL-VAE causes corruptions and blurs, Reducio-VAE generally maintains the subtle texture within the middle frame. Tab. 1 also reflects the advantage of Reducio-VAE in quantitative reconstruction metrics, e.g., PSNR, SSIM, and LPIPS on a randomly sampled validation subset of 1K videos with the duration of 1s on Pexels, and test on both UCF-101 and Pexels for FVD.
As shown in the table, Reducio-VAE strikes obvious advantages on overall reconstruction performance. Specifically, our model significantly outperforms state-of-the-art 2D VAE, e.g., SD2.1-VAE and SDXL-VAE, by more than 5 db in PSNR. Moreover, Reducio-VAE also performs better than VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2. Our VAE also outperforms the concurrent work, Cosmos-VAE, which also adopts a further compressed video latent space, by 0.2 in SSIM and 5 db in PSNR with an higher down-sampling factor.
We observe that Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence results in a lower reconstruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists of videos with much improved visual quality and much higher resolution.
Reducio-DiT is a powerful image-to-video model. As shown in Tab. 2, Reducio-DiT achieves a 90-point lower FVD in the zero-shot setting on UCF-101 compared to the state-of-the-art image-to-video model, DynamiCrafter. Additionally, Reducio-DiT surpasses VideoComposer in FVD metrics on MSR-VTT but falls short of DynamiCrafter in this aspect. For qualitative evaluation, we conduct a user study using identical images generated by PixArt--1024 as input. Participants evaluate 15 video groups based on three criteria: (1) visual quality, (2) motion quality, and (3) alignment of the condition image. Responses from 20 participants are presented in Fig. 6.
Notably, Reducio-DiT significantly outperforms in visual quality (64%) and content alignment (57%) by overcoming temporal flickering and preserving fine details from the condition image. Moreover, thanks to its strong temporal consistency, Reducio-DiT achieves a preference rate comparable to SVD-XT in motion quality, yet needs much fewer training hours and data. Fig. 6 displays the comparison between Reducio-DiT and the other two stat-of-the-art image-to-video models. While DynamicCrafter and SVD-XT fail to generate stable frames consistent with the given content frame, Reducio-DiT yields high visual quality with maintained facial details. Overall, Reducio-DiT demonstrates strong performance in image-to-video generation.
top: pouring milk into a cup of coffee. bottom: A storm trooper vacuuming the beach.
Reducio-DiT effectively balances efficiency and performance. By integrating Reducio-DiT with a text-to-image LDM for condition image generation, the factorized text-to-video task can be performed efficiently. Consequently, our model achieves FVD scores of 318.5 on UCF-101 and 291.0 on MSR-VTT and the VBench score of 81.39, beating a range of state-of-the-art video LDMs, as recorded in Tab. 3. More importantly, as most mainstream text-to-video models are built upon pre-trained text-to-image LDMs, we compare Reducio-DiT with baselines on training costs for video samples. Thanks to the extremely compressed latent space, we achieve competitive performance with a training cost of merely 3.2K A100 hours.
Moreover, Reducio-DiT significantly boosts the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output videos of 16 frames in almost 1 second, which realizes a speed-up compared with Lavie. Moreover, 16-frame video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speed-up of over Lavie. We argue that the core philosophy of Reducio lies in the extremely compressed latent space, which accelerates both the training and inference speed of text-to-video generation.
Although Reducio-DiT is trained exclusively on 16-frame videos sampled at 16 fps, we present that it can be adapted for longer video generation in a zero-shot manner. This is achieved by recursively using the last frame of a generated clip as the input condition for the subsequent clip. Fig. 7 presents the visualization of frames sampled uniformly from 32-frame and 48-frame videos, demonstrating that Reducio-DiT maintains consistent motion while preserving high visual quality.
4.3 Ablation Studies
4.3.1 Design of Reducio-VAE
| PSNR | SSIM | ||
| 16 | 16 | 34.14 | 0.93 |
| 2 | 32 | 36.79 | 0.96 |
| 4 | 32 | 35.88 | 0.94 |
| 1 | 64 | 35.51 | 0.93 |
Spatial redundancy is the key for compression. As Tab. 4 displays, when adopting an identical overall down-sampling factor , Reducio-VAE achieves worse performance with a larger compression factor in the temporal dimension. The result suggests that when a reference frame is provided, videos contain heavier redundancy in the spatial dimension than in the temporal dimension. Similarly, when is kept unchanged, increasing from 2 to 4 leads to a drop in PSNR by 0.91 db. To balance the trade-off between more compact latent space and better reconstruction performance, we set and in default.
Scaling up latent channel is helpful for improving the reconstruction performance, as results shown in Tab. 5. When grows from 4 to 16, the PSNR of reconstructed videos increases by 2.71 db. However, increasing the latent channel has a ceiling effect, as setting to 32 hardly achieves further gains in both PSNR and SSIM. Therefore, we opt to use latent space with for Reducio as the default setting. A visual qualitative comparison is presented in Fig. 8, where we compare the results when , , and .
| PSNR | SSIM | |
| 4 | 33.17 | 0.91 |
| 8 | 34.15 | 0.93 |
| 16 | 35.88 | 0.94 |
| 32 | 35.24 | 0.94 |
Fusion with cross-attention achieves the most competitive performance in Tab. 6. As for add-based fusion, we feed the content frame features into a convolution and replicate the output on the temporal dimension. Then we add video features with the content feature, together with a temporal embedding. Concerning linear-based fusion, we view the content condition as a single-frame video feature and concatenate it with the video feature by the temporal dimension. A 3D convolution layer then refines the concatenated features. As observed in Tab. 6, the attention-based Reducio-VAE outperforms other baselines by 1.71 in PSNR and 0.01 in SSIM, while demanding higher computational costs. By default, we display results using the attention-based fusion. Note that all Reducio-VAE displayed in Tab. 6 share the same 3D encoder and hence an identical latent space, where users are free to choose any of the decoders with different resource constraints.
| Fusion | PSNR | SSIM | Params |
| w/o. | 27.91 | 0.80 | 435M |
| linear | 30.86 | 0.87 | 439M |
| add | 34.17 | 0.93 | 450M |
| attention | 35.88 | 0.94 | 437M |
4.3.2 Design of Reducio-DiT
For ablation studies on DiT, we conduct training for 300 A100 hours for each experiment on Pexels and report FVD and IS scores on UCF-101.
Using patch size of one yields the best performance. As shown in Tab. 7, keeping a fixed spatial down-sampling factor of 32, we compare the performance of using a 2 higher temporal down-sampling factor in latent space and a temporal patch size of 1 (, ), with temporal patch size of 2 when patchify the input latents for obtaining patch embedding (, ). We observe that compressing the temporal dimension with VAE outperforms. Similarly, we ablate using VAE with 2 lower spatial down-sampling factor yet setting the patch size of spatial dimension to 2 (, ). The experimental results identify that compressing the video with Reducio-VAE to an extreme extent and using a small patch size, i.e., 1, in both the temporal and spatial dimensions achieves better performance.
| FVD | IS | ||||
| 2 | 32 | 2 | 1 | 351.0 | 32.7 |
| 4 | 16 | 1 | 2 | 534.1 | 30.4 |
| 4 | 32 | 1 | 1 | 337.6 | 34.1 |
A stunning sight as a steam train leaves the bridge, traveling over the arch-covered
viaduct. The landscape is dotted with lush greenery and rocky mountains.
Fusion with both semantic and content information contributes to the best quality. As shown in Tab. 8, using semantic features, i.e., OpenClip-based features alone may contribute to distorted visual detail, leading to a lower FVD. on the other hand, using content-based features helps to make video smooth and stable, achieving a higher FVD yet less versatile motion. A comparison example is illustrated in Fig. 9. The collaboration of content and semantic features helps Reducio strike the best balance between content consistency and motion richness.
| Condition | FVD | IS |
| semantic | 472.1 | 34.5 |
| content | 363.7 | 31.1 |
| semantic + content | 337.6 | 34.1 |
5 Conclusion
Video generation has shown promising results with many potential applications while it is still trapped by the unaffordable computation cost. In this paper, we explored how to effectively cut the overhead by reducing the latent size. In particular, we found that video can be encoded into extremely compressed latents with the help of a content image prior, where the latent codes only need to represent the motion variables. With this observation, we designed the Reducio-VAE to compress videos to smaller latents. Using this powerful Reducio-VAE, we trained the Reducio-DiT that enables fast high-resolution video generation. Our method is also compatible with other accelerating techniques, e.g., efficient attention, rectified flow, diffusion distillation, and etc., allowing for further speedup, where we leave it for future exploration.
Limitations and future work. While Reducio-DiT maintains strong consistency with the center frame, the number of training videos is still relatively limited for good instruction following performance. Also, 16-frame training samples somehow limit the magnitude of the motion in generated videos. As Reducio is much more compute-friendly with increased video frames and extended training data. Therefore, we believe in the potential of adapting our work for longer video generation on large-scale training and will explore it as future work.
Acknowledgement This work was supported in part by the National Natural Science Foundation of China (Grant 62032006 and Grant 62472098).
Appendix A Implementation details
Reducio-VAE. We demonstrate the detailed architecture of Reducio in Fig. 10. Generally, we follow the training strategies of SD-VAE [45]. We employ a customized version of PatchGAN [21] based on 3D convolutions and optimize the model with loss, KL loss, perceptual loss [82], and GAN loss. While initializing the 2D encoder and the 3D VAE with SD-VAE pre-trained weights accelerates convergence, we find that freezing the 2D encoder leads to worse performance than training the full parameters. We follow LaMD [19] to feed the motion latent into a normalization layer to obtain the output of the VAE encoder. Therefore, in the stage of diffusion training, we use a scale factor of 1.0, which is multiplied by the input latent as input into DiT.
| Reducio-VAE | ||
| -shape†††on videos | ||
| Channels (3d) | 128 | |
| Channels (2d) | 128 | |
| Ch Multiplier (3d) | 1,2,2,4,4,4 | |
| Ch Multiplier (2d) | 1,2,2,4,4 | |
| Depth | 2 | |
| Batch size | 32 | 24 |
| Learning rate | 4e-5 | |
| Iterations | 1,000,000 | |
During inference, We split videos with a resolution over into overlapping spatial tiles, we fuse the encoded latent as well as the video output, in a similar manner with Movie Gen [40]. Note that since Reducio-VAE employs a spatial down-sampling factor of 32, Reducio-VAE can only process video inputs whose width and height are divisible by 32. Otherwise, videos should be padded to meet this requirement before being fed into the VAE.
Reducio-DiT. We elaborate on the details of content image conditions in Fig. 11. During inference, we use classifier-free guidance [16] for better generation quality and set the default scale to 2.5. During training, we randomly drop image conditions at a probability of 0.1, as well as drop 10% text conditions.
Appendix B More ablations
| Conv | PSNR | SSIM | |
| 64 | 2d | 30.22 | 0.86 |
| 64 | 3d | 35.51 | 0.94 |
Using 3d VAE in Reducio-VAE helps to reconstruct videos in a better quality. We keep to 1 and to 32 and implement VAE with 2d convolutions. During decoding, we duplicate the middle frame condition for times to fuse with the latent of each frame respectively. As shown in Tab. 10, Reducio-VAE with 3d convolution outperforms its counterpart with 2d convolution. We believe that 3d convolution facilitates the VAE to model consistent motion and capture spatiotemporal differences.
| Content Frame | PSNR | SSIM |
| n/a | 27.91 | 0.80 |
| random | 31.72 | 0.87 |
| middle | 35.88 | 0.94 |
Using middle frame in Reducio-VAE. The content frame in Reducio-VAE provides a strong content prior and hence leads to a promising reconstruction performance. On the other hand, relying on any given frame as the content image may not generalize perfectly in all scenarios, especially when certain entities appear only briefly or outside the chosen frame. As shown in Tab. 11, we choose the middle frame by default as it serves as a more stable and robust content guidance due to its temporal centrality. Meanwhile, Reducio-VAE without condition achieves significantly worse results in both PSNR (-7.97) and SSIM (-0.14). In consequence, the Reducio-DiT framework without condition-based 3D VAE leads to unsatisfactory results featured with blurry frames and obvious visual defects.
| Model | latent shape | PSNR | SSIM | |
| DC-AE [7] | 32 | 30.68 | 0.70 | |
| Reducio-VAE | 16 | 35.56 | 0.97 |
Comparison between DC-AE and Reducio-VAE. We compare Reducio-VAE with DC-AE [7] on the Pexel test split with resolution of . As shown in the Table below, Reducio-VAE outperforms DC-AE on PSNR and SSIM by 4.88 and 0.27, respectively, highlighting the advantage of our framework in video domain.
| Attn | FVD | IS |
| 2d + 1d | 382.2 | 32.4 |
| 3d | 337.6 | 34.1 |
| Model | Quality | Semantic | Total |
| Score | Score | Score | |
| Show-1 [79] | 80.42 | 72.98 | 78.93 |
| Lavie [64] | 78.78 | 70.31 | 77.08 |
| VideoCrafter [5] | 81.59 | 72.22 | 79.72 |
| OpenSora v1.2 [85] | 81.35 | 73.39 | 79.76 |
| Lavie-2 [64] | 83.24 | 75.76 | 81.75 |
| Pyramid Flow [22] | 84.74 | 69.62 | 81.72 |
| VideoCrafter-2 [15] | 83.27 | 76.73 | 81.97 |
| Reducio-DiT | 82.24 | 78.00 | 81.39 |
| WAN [58] | 84.92 | 80.10 | 83.96 |
| STIV [26] | 81.20 | 72.70 | 79.50 |
| CausVid [75] | 85.21 | 78.57 | 83.88 |
| Model | subject | background | temporal | motion | dynamic | aesthetic | imaging | object | multiple | human | color | spatial | scene | appearance | temporal | overall |
| consistency | consistency | flickering | smoothness | degree | quality | quality | class | objects | action | relationship | style | style | consistency | |||
| Lavie [64] | 91.41 | 97.47 | 98.30 | 96.38 | 49.72 | 54.94 | 61.90 | 91.82 | 33.32 | 96.80 | 86.39 | 34.09 | 52.69 | 23.56 | 25.93 | 26.41 |
| Show-1 [79] | 95.53 | 98.02 | 99.12 | 98.24 | 44.44 | 57.35 | 58.66 | 93.07 | 45.47 | 95.60 | 86.35 | 53.50 | 47.03 | 23.06 | 25.28 | 27.46 |
| VideoCrafter [5] | 95.10 | 98.04 | 98.93 | 95.67 | 55.00 | 62.67 | 65.46 | 78.18 | 45.66 | 91.60 | 93.32 | 58.86 | 43.75 | 24.41 | 25.54 | 26.76 |
| OpenSora v1.2 [85] | 96.75 | 97.61 | 99.53 | 98.50 | 42.39 | 56.85 | 63.34 | 82.22 | 51.83 | 91.20 | 90.08 | 68.56 | 42.44 | 23.95 | 24.54 | 26.85 |
| Lavie-2 [64] | 97.90 | 98.45 | 98.76 | 98.42 | 31.11 | 67.62 | 70.39 | 97.52 | 64.88 | 96.40 | 91.65 | 38.68 | 49.59 | 25.09 | 25.24 | 27.39 |
| Pyramid Flow [22] | 96.95 | 98.06 | 99.49 | 99.12 | 64.63 | 63.26 | 65.01 | 86.67 | 50.71 | 85.60 | 82.87 | 59.53 | 43.20 | 20.91 | 23.09 | 26.23 |
| VideoCrafter-2 [15] | 97.17 | 98.54 | 98.46 | 97.75 | 42.50 | 65.89 | 70.45 | 93.39 | 49.83 | 95.00 | 94.41 | 64.88 | 51.82 | 24.32 | 25.17 | 27.57 |
| Reducio-DiT | 98.05 | 99.13 | 98.45 | 98.77 | 27.78 | 64.02 | 67.67 | 91.49 | 69.91 | 92.60 | 89.06 | 52.85 | 54.90 | 25.16 | 26.40 | 28.87 |
A dog wearing a Superhero outfit with red cape flying through the sky.
A dog wearing a Superhero outfit with red cape flying through the sky.
Using joint spatiotemporal 3D attention in Reducio-DiT outweighs using factorized spatial and temporal attention (i.e., 2D + 1D attention) in generation quality. Interestingly, we observe that factorized attention leads to a faster convergence of training loss. However, with the same training step, as shown in Tab. 13, factorized attention lags behind its counterpart with joint 3D attention for 45 in FVD. We suppose the possible reason is that 2D + 1D scheme demands adding additional temporal layers and performs factorized self-attention on a small set of tokens e ach, making it hard to model smooth open-set motion with the light computation. In contrast, 3D attention directly exploits the original parameters and collaborates all spatiotemporal tokens.
Quantitative results. We display the detailed performance comparison on VBench [20] in Tab. 15 and Tab. 14. Despite using only 3.2K A100 GPU hours and 5.4M training samples, Reducio-DiT achieves a promising semantic score of 78.00, beating a range of state-of-the-art LDMs. While the most recent models such as WAN [58] and CausVid [75] achieve higher overall scores than Reducio-DiT, we argue that our model uses a much smaller scale of training data and has a relatively small model scale, i.e., 1.2B.